# Breaking Annotation Barriers: Generalized Video Quality Assessment via Ranking-based Self-Supervision

### v2

@@ A.2 Analysis of Low-level Metrics
Our data selection strategy is based on a mixed-integer programming method [56], which optimizes dataset composition by aligning feature histograms. Specifically, we utilize this approach to match the distributions of nine low-level metrics (blockiness [45], blur [41], contrast [43], noise, flickering [42], colourfulness [14], luminance, spatial information (SI) [19], and temporal information (TI) [19]) between our dataset and widely used VQA datasets. Each metric is computed as follows:
@@
[45] is quantified by analyzing the luminance differences between pixels within and across encoding blocks. Specifically, we compute the absolute luminance differences between adjacent pixel pairs within the same encoding block (internal pixel pairs) and those spanning adjacent blocks (external pixel pairs). The blockiness metric is then determined as the ratio of the total sum of internal pixel difference values to the total sum of external pixel difference values across the entire video frame:
@@
is measured using the Cumulative Probability of Blur Detection (CPBD) [41], which evaluates perceptual sharpness based on edge width distribution. A higher CPBD value indicates a sharper image. Given an edge pixel $e_{i}$, its width $w(e_{i})$ is compared with the Just Noticeale Blur (JNB) threshold, determining the blur detection probability $w_{JNB}(e_{i})$. The final CPBD score is computed as:
@@
is a measure of the dispersion of pixel intensity values within the video frame and can be quantified using the standard deviation of grayscale intensities [43]. Specifically, for a grayscale image $I(x,y)$, the mean intensity $\mu$ is first computed as:
@@
occurs when an encoder skips macroblocks to conserve bitrate, especially in low-texture, slow-motion regions [42]. It is quantified by counting macroblock transitions from an “unupdated” to an “updated” state, with a threshold $T_{f}$ ensuring only significant changes are considered. The flickering metric is computed as:
@@
quantifies color distribution differences across RGB channels, following [14]. Given a frame with RGB channels $R,G,B$, we compute:
@@
A.3 More Details of PLVD-DS Datas
@@ A.3.1 Spatial Distortion
Table 5: An overview of our testing datasets
KoNViD-1k [16] 2017 1,200 1,200 540p 8 24, 25, 30 In-the-wild
LIVE-VQC [50] 2018 585 585 240p–1080p 10 30 In-the-wild
YouTube-UGC [60] 2019 1,380 1,380 360p–4K 20 30 In-the-wild
LSVQ [70] 2021 38,811 38,811 99p–4K 5–12 $<$ 60 In-the-wild
Waterloo-IVC-4K [26] 2019 1200 20 540p, 1080p, 4k 9-10 24, 25, 30 H.264 compression
LIVE-YT-HFR [33] 2021 480 16 1080p 6-10 24, 30, 60, 82, 98, 120 Frame rate, VP9 compression
LIVE-YT-Gaming [71] 2022 600 600 360p–1080p 8–9 30, 60 PGC, UGC
CGVDS [47] 2023 360 15 480p, 720p, 1080p 30 20, 30, 60 H.264 compression
KVQ [32] 2024 4200 600 - 3-8 - UGC
@@
A.3.2 Temporal Distortion
@@ A.4 More Details on Testing Datasets
Table 6: Performance comparison of our model in terms of prediction accuracy on four VQA datasets against advanced LLMs. Additionally, performance comparison of an ensemble of five methods versus a single model on the same datasets
[…]
LLaVA-NeXT-Video (8B) [74] 0.508 0.531 0.512 0.531 0.506
mPLUG-Owl3-8B [68] 0.547 0.534 0.605 0.548 0.543
InternVL2__\__5-8B [59] 0.566 0.629 0.583 0.596 0.588
Qwen2.5-7B-VL [2] 0.696 0.666 0.740 0.664 0.631
LLaVA-ov-chat [24] 0.667 0.641 0.727 0.664 0.639
MinimalisticVQA(VII) [52] 0.907 0.817 0.906 0.839 0.784
MinimalisticVQA(IX)  [52] 0.926 0.845 0.904 0.843 0.820
FAST-VQA [61] 0.918 0.836 0.918 0.832 0.765
DOVER [64] 0.922 0.840 0.915 0.860 0.812
Q-Align [65] 0.922 0.849 0.923 0.846 0.822
Ensemble five methods 0.939 0.864 0.935 0.860 0.831
Soft Ranking (Stage 1) 0.926 0.855 0.935 0.849 0.853
Appendix B More Details of Pairwise Quality Annotation
B.1 VQA Models for Pseudo-labeling
We choose five SOTA VQA models: Minimalistic-VQA (VII) [52], Minimalistic-VQA (IX) [52], FAST-VQA [61], DOVER [64], and Q-Align [65] as our initial judges to formulate our pairwise quality annotation. As shown in Table 6, the ranking accuracy of the five ensemble methods is higher than that of any single model. The detail introduction of the five models is as follows:
Minimalistic-VQA (VII)
employs Swin Transformer-B [31], pre-trained on ImageNet-1K [9], as the spatial quality analyzer to extract quality-aware spatial features from key frames, ensuring robust spatial quality assessment.
Minimalistic-VQA (IX)
builds upon Minimalistic-VQA (VII) by incorporating a temporal quality analyzer to account for motion distortions. The temporal quality analyzer, implemented using the SlowFast [12] network pre-trained on the Kinetics-400 [5] dataset, extracts motion-related features from video chunks, enhancing the model’s ability to assess temporal quality variations.
@@ B.2 Prompts for Model Training
We final construct the label prompt of our large dataset using a fixed template:
Question: "Now you will receive two videos. The first video:<image>. The second video:<image>. Please watch these videos carefully, and then answer the following question: Comparing with the first video, how do you assess the quality of the second video?"
Answer: "The quality of the second video is [level] to the first video."
Appendix C More Details of Our Model
C.1 Training Details
The model is trained using the DeepSpeed framework with mixed-precision floating-point operations to optimize memory and computational efficiency. The training is conducted for one epoch with a batch size of 1 per device and a gradient accumulation step of 1. The optimizer follows AdamW with a learning rate of $1\times 10^{-5}$, a cosine learning rate schedule, and a warmup ratio of 0.03.
@@ C.1 Training Details
C.2 Inferring Details
C.2.1 Probability Modeling
@@
First, we describe how to construct the probability distribution for comparative quality assessments. For hard ranking, the comparative token set is defined as:
$\mathcal{S}=\{s_{k}\}_{k=1}^{2}=\{\textit{worse},\textit{better}\}.$ (21)
For soft ranking, the comparative token set is extended to:
@@
where $\alpha_{k}$ are fixed weights that reflect the comparative levels. Specifically, for hard ranking, the weights are:
$\{\alpha_{k}\}_{k=1}^{2}=\{0,0.5\}.$ (25)
For soft ranking, the weights are defined as:
@@ C.2.2 Score Modeling
With this probability matrix, we estimate the final quality score using maximum a posteriori (MAP) [54] estimation under Thurstone’s Case V model [53]. This is formulated as the following convex optimization problem:
@@ Appendix D More Details of Experimental Results
We also compare the prediction accuracy of our model with several advanced LLMs, including LLaVA-NeXT-Video (7B) [74], mPLUG-Owl3-8B [68], InternVL2.5-8B [59], Qwen2.5-7B-VL [2], and our base model LLaVA-ov-chat [24], as shown in Table 6. Evidently, our model significantly outperforms all LLMs, suggesting that existing LLMs trained on high-level tasks still struggle with low-level visual perception.

### v3

@@ A.2 Analysis of Low-level Metrics
Our data selection strategy is based on a mixed-integer programming method (Vonikakis et al., 2017), which optimizes dataset composition by aligning feature histograms. Specifically, we utilize this approach to match the distributions of nine low-level metrics (blockiness (Romaniak et al., 2012), blur (Narvekar & Karam, 2011), contrast (Peli, 1990), noise, flickering (Pandel, 2008), colourfulness (Hasler & Suesstrunk, 2003), luminance, spatial information (SI)  (ITU-T P.910, 2008), and temporal information (TI) (ITU-T P.910, 2008)) between our dataset and the LSVQ dataset. Each metric is computed as follows:
Figure 9: Distribution of nine metrics on the LSVQ dataset, as well as on our dataset before and after sampling.
@@
(Romaniak et al., 2012) is quantified by analyzing the luminance differences between pixels within and across encoding blocks. Specifically, we compute the absolute luminance differences between adjacent pixel pairs within the same encoding block (internal pixel pairs) and those spanning adjacent blocks (external pixel pairs). The blockiness metric is then determined as the ratio of the total sum of internal pixel difference values to the total sum of external pixel difference values across the entire video frame:
@@
is measured using the Cumulative Probability of Blur Detection (CPBD) (Narvekar & Karam, 2011), which evaluates perceptual sharpness based on edge width distribution. A higher CPBD value indicates a sharper image. Given an edge pixel $e_{i}$, its width $w(e_{i})$ is compared with the Just Noticeale Blur (JNB) threshold, determining the blur detection probability $w_{JNB}(e_{i})$. The final CPBD score is computed as:
@@
is a measure of the dispersion of pixel intensity values within the video frame and can be quantified using the standard deviation of grayscale intensities (Peli, 1990). Specifically, for a grayscale image $I(x,y)$, the mean intensity $\mu$ is first computed as:
@@
occurs when an encoder skips macroblocks to conserve bitrate, especially in low-texture, slow-motion regions (Pandel, 2008). It is quantified by counting macroblock transitions from an “unupdated” to an “updated” state, with a threshold $T_{f}$ ensuring only significant changes are considered. The flickering metric is computed as:
@@
quantifies color distribution differences across RGB channels, following (Hasler & Suesstrunk, 2003). Given a frame with RGB channels $R,G,B$, we compute:
@@
A.3 More Details on Synthetic Distortion Data
@@ A.3.1 Spatial Distortions
Table 3: An overview of our testing datasets.
KoNViD-1k (Hosu et al., 2017) 2017 1,200 1,200 540p 8 24, 25, 30 In-the-wild
LIVE-VQC (Sinno & Bovik, 2018) 2018 585 585 240p–1080p 10 30 In-the-wild
YouTube-UGC (Wang et al., 2019) 2019 1,380 1,380 360p–4K 20 30 In-the-wild
LSVQ (Ying et al., 2021) 2021 38,811 38,811 99p–4K 5–12 << 60 In-the-wild
Waterloo-IVC-4K (Li et al., 2019b) 2019 1200 20 540p, 1080p, 4k 9-10 24, 25, 30 H.264 compression
LIVE-YT-HFR (Madhusudana et al., 2021) 2021 480 16 1080p 6-10 24, 30, 60, 82, 98, 120 Frame rate, VP9 compression
LIVE-YT-Gaming (Yu et al., 2022) 2022 600 600 360p–1080p 8–9 30, 60 PGC, UGC
CGVDS (Saha et al., 2023) 2023 360 15 480p, 720p, 1080p 30 20, 30, 60 H.264 compression
KVQ (Lu et al., 2024) 2024 4200 600 - 3-8 - UGC
A.3.2 Temporal Distortions
@@ A.4 More Details on Testing Datasets
Appendix B More Details of Quality Annotation
B.1 Weak Models for Pseudo-labeling
Table 4: Comparison of model parameters and architecture.
Model Parameters (M) Architecture
MinimalisticVQA(VII) 86.93 Swin-B
MinimalisticVQA (IX) 121.59 Swin-B + SlowFast
FAST-VQA 29.97 Swin-Tiny
DOVER 58.06 Swin-Tiny + Conv-Tiny
Q-Align 8204.56 mPLUG-Owl2
Our strong model 8075.24 LLaVA-OneVision-Chat + SlowFast
We choose five SOTA VQA models: MinimalisticVQA (VII) (Sun et al., 2024), MinimalisticVQA (IX) (Sun et al., 2024), FAST-VQA (Wu et al., 2022), DOVER (Wu et al., 2023a), and Q-Align (Wu et al., 2023b) as weak teachers to formulate our pseudo quality annotation. The detail introduction of the five models is as follows:
MinimalisticVQA (VII)
employs Swin Transformer-B (Liu et al., 2022), pre-trained on ImageNet-1K (Deng et al., 2009), as the spatial quality analyzer to extract quality-aware spatial features from key frames, ensuring robust spatial quality assessment.
MinimalisticVQA (IX)
builds upon MinimalisticVQA (VII) by incorporating a temporal quality analyzer to account for motion distortions. The temporal quality analyzer, implemented using the SlowFast (Feichtenhofer et al., 2019) network pre-trained on the Kinetics-400 (Carreira & Zisserman, 2017) dataset, extracts motion-related features from video chunks, enhancing the model’s ability to assess temporal quality variations.
@@ B.2 Prompts for Model Training
We construct the label prompts for our large-scale dataset using a fixed template. For the single-video input:
Question: "You will now receive a video: <image>. Please watch the video carefully and answer the following question: What is your overall rating of the quality of this video?"
Answer: "[quality score]"
For the dual-video input:
Question: "You will now receive two videos. The first video: <image>. The second video: <image>. Please watch both videos carefully and answer the following question: Compared to the first video, how would you rate the quality of the second video?"
Answer: "The quality of the second video is [level] compared to the first video."
Appendix C More Details of Our Strong student Model
C.1 Model Structure
As illustrated in Fig. 3, our model comprises three components: a visual feature extractor, a text tokenizer, and an LLM decoder.
$\bm{r}$ $=\mathcal{L}(\bm{f}^{v}_{A},\bm{f}^{v}_{B},\bm{f}^{p}),$ (22)
where $\bm{r}$ is expected to belong to {“superior”, “better”, “similar”, “worse”, “inferior”}.
C.2 Training Details
C.2.1 Training Setup
The model is trained using the DeepSpeed framework with mixed-precision floating-point operations to optimize memory and computational efficiency. The training is conducted for one epoch with a batch size of 2 per device and a gradient accumulation step of 1. The optimizer follows AdamW with a initial learning rate of $1\times 10^{-4}$, a cosine learning rate schedule, and a warm-up ratio of 0.03.
@@ C.2.1 Training Setup
C.2.2 Auxiliary Confidence Loss
As mentioned in Section 4.3, we introduce an auxiliary confidence loss to encourage the model to maintain high-confidence predictions, especially in the presence of noisy weak supervision. The final training objective is a dynamically weighted combination of the cross-entropy loss $\mathcal{L}_{\text{CE}}$ and the confidence loss $\mathcal{L}_{\text{conf}}$:
$\mathcal{L}=(1-\lambda)\cdot\mathcal{L}_{\text{CE}}+\lambda\cdot\mathcal{L}_{\text{conf}},$ (23)
where $\lambda$ is an adaptive weighting factor that balances between trusting the weak labels and relying on the model’s own confidence. The confidence loss is defined as the average entropy over the predicted token probability distributions:
$\mathcal{L}_{\text{conf}}=\frac{1}{N}\sum_{i=1}^{N}H(p_{\theta}(x_{i}))=-\frac{1}{N}\sum_{i=1}^{N}\sum_{c}p_{\theta}(c|x_{i})\log p_{\theta}(c|x_{i}),$ (24)
where $p_{\theta}(c|x_{i})$ denotes the predicted probability of vocabulary token $c$ given input $x_{i}$. By minimizing the entropy of the predicted distribution, we encourage the model to produce more confident next-token predictions.
To dynamically adjust $\lambda$ during training, we introduce a temperature-based confidence estimation mechanism. Specifically, we define:
$\lambda=\alpha\cdot\min\left(1.0,\frac{t}{T_{\text{warmup}}}\right),$ (25)
where $t$ denotes the current training step ratio (normalized to $[0,1]$), and $T_{\text{warmup}}$ is the warm-up period, which we set to $10\%$ of the total training steps. This warm-up phase ensures that the strong model gradually learns to rely on its own confidence, while initially being guided by the weak labels. The factor $\alpha$ is computed as the ratio between the temperature-scaled exponentials of the two losses:
$\alpha=\frac{\exp(\mathcal{L}_{\text{conf}}/T)}{\exp(\mathcal{L}_{\text{conf}}/T)+\exp(\mathcal{L}_{\text{CE}}/T)}.$ (26)
Here, $T$ is a temperature parameter that controls the sharpness of the weighting between the two loss components. We linearly decrease $T$ from $0.5$ to $0.1$ during the warm-up period to gradually increase the sensitivity of $\alpha$ to differences in the two loss values.
C.3 Inferring Details
C.3.1 Probability Modeling
@@
First, we describe how to construct the probability distribution for comparative quality assessments. The comparative token set is defined as:
@@
where $\alpha_{k}$ are fixed weights that reflect the comparative levels. Specifically, the weights are defined as:
@@ C.3.2 Score Modeling
Table 5: Performance of weak-to-strong models trained with pseudo-labels from weak models. For comparison, we also report the performance of our model trained directly on the LSVQ dataset.
In-domain Datasets $_{\text{test}} LSVQ$$_{\text{1080p}}$ KoNViD-1k LIVE-VQC YouTube-UGC Overall
# of videos 7,182 3,573 1,200 585 1,020 -
Methods SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
Weak Teachers
MinimalisticVQA(VII) 0.861 0.859 0.740 0.784 0.843 0.841 0.757 0.813 0.775 0.779 0.817 0.830
MinimalisticVQA(IX) 0.885 0.882 0.792 0.828 0.862 0.859 0.775 0.821 0.826 0.821 0.849 0.859
FAST-VQA 0.880 0.880 0.781 0.813 0.859 0.854 0.826 0.845 0.730 0.747 0.838 0.849
DOVER 0.878 0.866 0.782 0.813 0.874 0.869 0.817 0.840 0.771 0.781 0.842 0.845
Q-Align 0.886 0.884 0.761 0.822 0.876 0.878 0.783 0.819 0.834 0.846 0.844 0.861
Weak-to-Strong Students
MinimalisticVQA(VII)-labeled 0.855 0.852 0.762 0.795 0.859 0.857 0.771 0.813 0.808 0.821 0.824 0.833
MinimalisticVQA(IX)-labeled 0.879 0.878 0.794 0.826 0.869 0.871 0.786 0.822 0.843 0.846 0.849 0.859
FAST-VQA-labeled 0.871 0.868 0.785 0.819 0.849 0.855 0.798 0.833 0.825 0.834 0.840 0.850
DOVER-labeled 0.877 0.869 0.780 0.813 0.870 0.875 0.792 0.829 0.819 0.831 0.843 0.850
Q-Align-labeled 0.878 0.876 0.794 0.824 0.873 0.880 0.781 0.825 0.833 0.853 0.848 0.859
Supervised Student
LSVQ-labeled 0.881 0.878 0.797 0.834 0.874 0.874 0.797 0.828 0.830 0.838 0.851 0.861
# of videos 600 357 480 1,200 2,926 -
Methods SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
Weak Teachers
MinimalisticVQA(VII) 0.596 0.682 0.681 0.733 0.061 0.130 0.275 0.338 0.604 0.659 0.490 0.551
MinimalisticVQA(IX) 0.686 0.746 0.797 0.816 0.301 0.388 0.459 0.502 0.615 0.661 0.574 0.622
FAST-VQA 0.631 0.677 0.725 0.747 0.326 0.415 0.327 0.363 0.518 0.526 0.486 0.512
DOVER 0.647 0.728 0.694 0.747 0.360 0.465 0.368 0.418 0.559 0.593 0.519 0.569
Q-Align 0.611 0.681 0.756 0.798 0.329 0.342 0.414 0.497 0.613 0.655 0.555 0.606
Weak-to-Strong Students
MinimalisticVQA(VII)-labeled 0.632 0.717 0.718 0.773 0.318 0.386 0.356 0.412 0.604 0.652 0.536 0.593
MinimalisticVQA(IX)-labeled 0.687 0.748 0.763 0.810 0.383 0.461 0.459 0.515 0.638 0.676 0.591 0.639
FAST-VQA-labeled 0.658 0.766 0.752 0.785 0.392 0.422 0.414 0.493 0.585 0.624 0.550 0.604
DOVER-labeled 0.662 0.758 0.752 0.809 0.449 0.482 0.435 0.519 0.574 0.627 0.554 0.617
Q-Align-labeled 0.671 0.738 0.744 0.785 0.437 0.480 0.450 0.525 0.620 0.668 0.581 0.636
Supervised Student
LSVQ-labeled 0.643 0.713 0.713 0.770 0.451 0.490 0.451 0.485 0.619 0.636 0.577 0.608
@@
With this probability matrix, we estimate the final quality score using maximum a posteriori (MAP) (Tsukida et al., 2011) estimation under Thurstone’s Case V model (Thurstone, 2017). This is formulated as the following convex optimization problem:
@@ Appendix D More Details of Experimental Results
D.1 More Details of Weak-to-strong Generalization Effect
Table 5 presents the per-dataset results from the experiments described in Section 3.3. For in-domain benchmarks, the student model achieves performance comparable to its teachers, with slight improvements, demonstrating that our simple knowledge distillation approach effectively transfers quality assessment knowledge from weak to strong models. For OOD benchmarks, the student model shows substantial improvements over its teachers, highlighting a pronounced weak-to-strong generalization effect.